OpenStudy (fluttershyk):

A bag contains 2 white marbles and 7 purple marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second is drawn. A. What is the probability of selecting a purple marble and then a white marble? B. What is the probability of selecting two white marbles? C. Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row? Show your Work.

OpenStudy (fluttershyk):

@rebeccaxhawaii @ganeshie8 @KamiBug @Cas_fangirl_14 @SuperGirl22

OpenStudy (fluttershyk):

@supersmart1001

OpenStudy (gm100299):

a 2+7 = 9 (the total number of marbles) 7/9 * 2/8 = 14/72 and reduce

OpenStudy (gm100299):

b 2/9 * 1/8 = 2/72 and reduce

OpenStudy (gm100299):

c since (b) already shows the probability of choosing 2 white in a row you just need to work out the purple. 7/9 * 6/8 = 42/72 and reduce

OpenStudy (fluttershyk):

@GM100299 thanks but I already figured it out but ill still medal and fan you